# StdGEN: Semantic-Decomposed 3D Character Generation from Single Images

### v1

@@ 1 Introduction
To address these limitations, CharacterGen [40] was developed to efficiently generate characters from single images using a multi-view diffusion model and large reconstruction model [17]. Despite showing impressive generation capabilities in various posed images, CharacterGen can only produce composed avatars in watertight meshes with no decomposability. These meshes require significant manual labor to separate, edit, or animate, limiting their applicability. Moreover, the quality of the generated meshes is often unsatisfactory, particularly in finer details such as the character’s face and clothing, as shown in Fig. 5. Therefore, efficiently generating high-quality, decomposable 3D characters remains an open challenge.
@@
  • Building upon (1) our S-LRM, (2) an efficient multi-view diffusion model and (3) iterative multi-layer mesh refinement, we present StdGEN, an efficient pipeline for high-quality decomposable 3D character generation, which for the first time enables semantic-decomposed 3D character creation from single images with arbitrary poses in just a few minutes.
@@ 2.2 Large Reconstruction Model
3D character generation is a challenging problem due to its high precision requirements and the scarcity of data. One line of work leverages 3D-aware GANs to model the distribution of digital humans  [2, 15, 24, 69, 37]. Recently the SDS-based methods have shown the possibility of generating a variety of stylized characters [3, 22, 59, 12], yet it suffers from the long optimization times and the difficulty of meticulous style control. The most related work to us is CharaterGen [40], which calibrates input poses to canonical multi-view images via an image-conditioned multi-view diffusion model, followed by LRM for 3D character reconstruction and multi-view texture back projection; however, it still exhibits limited geometry and texture quality. Furthermore, the methods above treat the parts of the character as a whole, requiring extra manual labor to process. Our approach, in contrast, employs a semantic-aware, feed-forward paradigm that generates high-quality, decomposable characters using only one forward pass, providing significant efficiency and quality improvement.
@@ 3.1 Multi-view Generation and Canonicalization
Multi-view RGBs and Normals Generation. This stage aims to provide more comprehensive information from different camera views for the subsequent semantic-aware large reconstruction model. Additionally, multi-view consistent normals are crucial to capturing rich surface details for further mesh refinement, which cannot be easily achieved by independently predicting normals for each view. Considering these factors, we adapt Era3D [26] to simultaneously generate multi-view high-resolution RGBs and normals. By employing memory-efficient row-wise attention mechanisms between different views and between RGB and normal maps, we ensure fine-grained spatial consistency and unlock the potential for higher output resolutions. Specifically, it accepts a single A-pose image as input and generates orthographic projections for six viewpoint RGBs and normals with the elevation of $0^{\circ}$ and azimuths of $\{-90^{\circ},-45^{\circ},0^{\circ},45^{\circ},90^{\circ},180^{\circ}\}$. To better suit our task and enhance the clarity of character details, we increased the resolution through a progressive training approach, ultimately reaching 1024.
@@ 3.2 Semantic-aware Large Reconstruction Model
Once obtaining multi-view images, [66, 40] use a transformer-based sparse-view Large Reconstruction Model (LRM) to reconstruct a holistic 3D mesh. In contrast to them, we aim to generate the character with the decomposed components, including the minimal-clothed human model, external clothing, and hair, to produce 3D character models that are not only visually accurate but also functionally versatile for various uses in 3D games and animation pipelines. To achieve this goal, we proposed the Semantic-aware Large Reconstruction Model (S-LRM), which extends NeRF/SDF to simultaneously encode semantics, appearance and geometry information in a feed-forward manner.
As shown in Fig. 2, the S-LRM takes multiview images into tokens and then feeds into a transformer-based image-to-triplane decoder to output a triplane representation. The triplane features are subsequently decoded to obtain semantic, color, density/SDF information. To enhance the training efficiency and reconstruction quality, following InstantMesh [66], we first utilize triplane NeRF representation to render 2D images with volume rendering for training. After that, we then switch to FlexiCubes [51] to extract explicit mesh from triplane-decoded SDF grid, and use direct mesh rasterization to render images. However, to obtain semantic-decomposed surface reconstruction, both NeRF and SDF implicit representations must be capable of rendering distinct semantic layers into images or extracting separate semantic surfaces using FlexiCubes in a differentiable manner. To achieve that, a novel semantic-equivalent NeRF/ SDF is proposed to extract character parts by specific semantics.
@@
If the probability of a certain semantic at a given location is zero, it should be considered completely transparent under the current semantic category. Furthermore, given that a position is known to be opaque, the probability of the current semantic should be linear to the final equivalent transparency.
@@
Model Structure. Our S-LRM’s core structure is derived from InstantMesh [66] with a ViT encoder, an image-to-triplane transformer, and the feature decoder. Several enhancements are made to this foundation to better suit our objectives, with the semantic decoder being a key addition. This component mirrors the structure of the density and color decoders and is designed to generate the semantic field in 3D space. For memory efficient training, we incorporate LoRA [19, 42] structures into all linear layers inside both self-attention and cross-attention blocks.
Multiple semantic level supervision. Current LRMs typically rely solely on 2D supervision, which limits their ability to generate information about objects’ internal structures under occlusion. While 3D supervision would be effective, it is often too resource-intensive. To address this, we propose an effective supervision method that jointly learns semantic and color information, enabling the acquisition of a 3D semantic field and internal character information using only 2D supervision.
Stage 1: Training on NeRF with Single-layer Semantics. In this initial stage, we train on the triplane NeRF representation. We initialize the model with the pre-trained InstantNeRF, and train the newly added LoRA in all attention blocks’ linear layers, and the newly introduced semantic decoder. We train it under the image, mask and semantic loss:
@@
where $\mbox{where }\ T_{P,i}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}\sum_{s\in P}p_{s,i}),$ (8)
The loss function is defined as:
Stage 3: Training on Mesh with Multi-layer Semantics. We switch to mesh representation [51] similar to [66] for efficient high-resolution training. We then extract the equivalent SDF via:
$f_{i,P}=\max(f_{i},(\max_{s\notin P}p_{i,s}-\max_{s\in P}p_{i,s})),$ (10)
$\mathcal{L}_{3}$ $=\mathcal{L}_{2}+\lambda_{\text{normal}}\sum_{i}M^{gt}_{P}\otimes\left(1-\hat{N}_{P,i}\cdot N_{P,i}^{gt}\right)$
$+\lambda_{\text{depth}}\sum_{i}M^{gt}_{P}\otimes\left\|\hat{D}_{P,i}-D_{P,i}^{gt}\right\|_{1}+\lambda_{\text{dev}}\mathcal{L}_{\text{dev}}$ (11)
where $\hat{D}_{P,i}$, $D_{P,i}^{gt}$, $\hat{N}_{P,i}$, $N_{P,i}^{gt}$ and $M_{P}^{gt}$ denote the rendered depth, ground truth depth, rendered normal, ground truth normal and ground truth mask of the $k$-th view under semantic set $P$, respectively; $\mathcal{L}_{\text{dev}}$ denotes the deviation loss of FlexiCubes.
3.3 Multi-layer Refinement
Given the limitations in geometric and texture detail achievable by large reconstruction models, further optimization of the mesh post-reconstruction is necessary. Recent methods [65, 27] utilizing high-resolution normal maps for original mesh optimization have shown promising results, albeit primarily for single-layer meshes. We propose an iterative optimization mechanism for multi-layer mesh refinement.
To ensure thorough optimization at each level, we employ a staged approach following our S-LRM: Initially, we extract different parts of the mesh by specifying various semantics and optimize only the base human model; upon completion, we overlay the clothing mesh, fixing the base and optimizing solely the clothing component; finally, we add the hair mesh, fixing the previous two layers and optimizing only the hair. The optimization process is guided by the multi-view normal maps generated earlier through diffusion models, each optimization step involves a differentiable rendering of the mesh to compute losses and gradients, followed by vertex adjustments based on these gradients, and re-mesh operations including edge collapse, split and flips. The loss function for this process is defined as follows:
Pre- and Post-dilation of Mesh. To better solve the problematic intersections among meshes, we introduce a ”dilation” process applied both before and after optimization. This process constructs an approximate ”flow field” based on the original positions of the inner and outer layer meshes.
4 Experiments
@@
2D Multi-view Comparisons
SyncDreamer [31] 0.870 0.183 0.223 0.864 0.845 0.217 0.328 0.839
@@
3D Character Comparisons
Magic123 [43] 0.886 0.142 0.192 0.887 0.849 0.197 0.256 0.862
@@
Table 1: Quantitative comparison of A-pose and arbitrary pose inputs for 2D multi-view generation and 3D character generation on the test split of Anime3D++ dataset.
@@ 4.1 Anime3D++ Dataset
While the Anime3D Dataset from CharacterGen [40] comprises RGB renderings of 13,746 stylized character subjects, it falls short of high-quality, decomposable character generation requirements. To address this limitation, we have developed the Anime3D++ Dataset. Following previous approaches [40, 7], we initially collected approximately 14,000 3D anime character models from VRoid-Hub. After a rigorous quality control process, including removing low-quality data and models unsuitable for proper layering, we refined our dataset to 10,811 character models. All characters were adjusted to an A-pose configuration, with arms rotated 45 degrees downward from the horizontal position.
Our rendering process goes beyond standard image generation, incorporating depth, normal, and semantic maps. To facilitate layered reconstruction supervision, we rendered the complete model and two additional configurations: the base human model alone, and the base model with clothing. The rendering includes eight views at 45-degree azimuth intervals with zero elevation, supplemented by top-down and bottom-up views. We enriched the dataset with five close-up facial views and 20 random viewpoints. To enhance the training of our diffusion model, we implemented data augmentation on varying arm angles.
4.2 Implementation Details
We divide our Anime3D dataset into a training and testing set in a 99:1 ratio. We first train the canonicalization diffusion model at a 512 resolution with a learning rate of 5e-5, then reduce it to 1e-5 as we progressively increase the resolution to 768 and 1024. Similarly, the multi-view diffusion model is trained at a constant learning rate of 5e-5 while scaling from 512 to 768 and finally to 1024 resolution. We use LoRA with a 128-rank for S-LRM, a learning rate of 4e-5, and three supervision stages with rendering resolutions of 192, 144, and 512. The loss function parameters are set as $\lambda_{\text{lpips}},\lambda_{\text{mask}},\lambda_{\text{sem}},\lambda_{\text{depth}},\lambda_{\text{normal}},\lambda_{\text{dev}}=2.0,1.0,1.0,0.5,0.2,0.5$. For multi-layer refinement, we set $\lambda_{\text{mask}}^{\prime},\lambda_{\text{normal}}^{\prime},\lambda_{\text{col}}=1.0,1.0,100.0$, and we further extract the coarse hair mask, applying additional normal and mask loss for hair refinement with a weight of 1 and 10.
@@ 4.3 Results and Comparisons
User Study. To comprehensively evaluate our method’s performance, we conducted a user study involving 28 volunteers. The study utilized 16 randomly selected 3D meshes and their corresponding reference images, drawn from both web-collected images and the Anime3D++ test split. Participants were asked to compare our method against existing approaches across four dimensions: overall quality, fidelity, geometric quality, and texture quality. For each comparison group, participants identified the best result in each dimension. Tab. 2 shows that our method achieved superior performance across all evaluated dimensions, demonstrating its effectiveness in generating high-quality 3D characters.
@@
4.4 Ablation Study
@@
Multi-layer Refinement. We further illustrate the distinction between the direct output of our S-LRM and the results after multi-layer refinement in Fig. 8. The pre-optimization results demonstrate that our S-LRM successfully decomposes various mesh components with correct geometry and shape, validating the capabilities of our approach. However, the precision is limited due to the inherent characteristics of FlexiCubes and memory constraints. Post-refinement, we observe a substantial enhancement in precision while maintaining the overall structure. This improvement underscores the effectiveness of our multi-layer refinement process in preserving the structure of the decomposed components while significantly elevating the geometric accuracy and overall quality of the reconstructed character.
4.5 Applications

### v2

@@ 1 Introduction
To address these limitations, CharacterGen [39] was developed to efficiently generate characters from single images using a multi-view diffusion model and large reconstruction model [15]. Despite showing impressive generation capabilities in various posed images, CharacterGen can only produce holistic avatars in watertight meshes with no decomposability. These meshes require significant manual labor to separate, edit, or animate, limiting their applicability. Moreover, the quality of the generated meshes is often unsatisfactory, particularly in finer details such as the character’s face and clothing, as shown in Fig. 5. Therefore, efficiently generating high-quality, decomposable 3D characters remains an open challenge.
@@
  • Building upon (1) our S-LRM, (2) an efficient multi-view diffusion model and (3) iterative multi-layer mesh refinement, we present StdGEN, an efficient pipeline for high-quality decomposed 3D character generation, which for the first time enables semantic-decomposed 3D character creation from arbitrary-posed single images in minutes.
@@ 2.2 Large Reconstruction Model
3D character generation is a challenging problem due to its high precision requirements and the scarcity of data. One line of work leverages 3D-aware GANs to model the distribution of digital humans  [2, 14, 21, 71, 36]. Recently the SDS-based methods have shown the possibility of generating a variety of stylized characters [3, 19, 56, 11, 22], yet it suffers from the long optimization times and the difficulty of meticulous style control. Frankenstein [65] concentrates on producing decomposed, textureless 3D meshes based on 2D layouts, restricting the potential for achieving high-fidelity reconstruction from the reference image. CharacterGen [39] calibrates input poses to canonical multi-view images via an image-conditioned multi-view diffusion model, followed by LRM for 3D character reconstruction and multi-view texture back projection, but still exhibits limited geometry and texture quality. Our approach, in contrast, employs a semantic-aware, feed-forward paradigm that generates high-quality, decomposable characters using only one forward pass from an arbitrary reference image, providing significant efficiency and quality improvement.
@@ 3.1 Multi-view Generation and Canonicalization
Multi-view RGBs and Normals Generation. This stage provides comprehensive information for the subsequent S-LRM, with multi-view consistent normals crucial for capturing rich surface details in mesh refinement rather than relying on independent normal predictions. We adapt Era3D [26] to generate high-resolution RGBs and normals simultaneously. Using memory-efficient row-wise attention across views and between RGB and normal maps, we ensure fine-grained spatial consistency and enable higher output resolutions. Specifically, it takes a single A-pose image as input, producing orthographic projections of six viewpoints (elevation $0^{\circ}$, azimuth ${-90^{\circ},-45^{\circ},0^{\circ},45^{\circ},90^{\circ},180^{\circ}}$). For sharper character details, we increased the resolution through a progressive training approach, ultimately reaching 1024.
@@ 3.2 Semantic-aware Large Reconstruction Model
Once obtaining multi-view images, [63, 39] use transformer based sparse-view Large Reconstruction Model (LRM) to reconstruct a holistic 3D mesh. In contrast, we aim to generate characters with decomposed components, including the minimal-clothed human model, external clothing, and hair, to produce 3D character models that are not only visually accurate but also functionally versatile for various uses in 3D games and animation pipelines. To achieve this goal, we proposed the Semantic-aware Large Reconstruction Model (S-LRM), which extends NeRF/SDF to simultaneously encode semantics, appearance and geometry information in a feed-forward manner.
As shown in Fig. 2, the S-LRM takes multi-view images into tokens and then feeds into a transformer-based image-to-triplane decoder to output a triplane representation. The triplane features are decoded to obtain semantic, color, density/SDF information. To enhance the training efficiency and reconstruction quality, following InstantMesh [63], we first utilize triplane NeRF representation to render 2D images with volume rendering for training. We then switch to FlexiCubes [49] to extract explicit mesh from triplane-decoded SDF grid, and use direct mesh rasterization to render images. However, to obtain semantic-decomposed surface reconstruction, both NeRF and SDF implicit representations must be capable of rendering distinct semantic layers into images or extracting separate semantic surfaces using FlexiCubes in a differentiable manner. To achieve that, a novel semantic-equivalent NeRF/ SDF is proposed to extract character parts by specific semantics.
@@
If the probability of a certain semantic at a given location is zero, it should be considered fully transparent under the current semantic category. Furthermore, given that a position is known to be opaque, the probability of the current semantic should be linear to the final equivalent transparency.
@@
Model Structure. Our S-LRM’s core structure is derived from InstantMesh [63] with a ViT encoder, an image-to-triplane transformer, and the feature decoder. Several enhancements are made to this foundation to better suit our objectives, with the semantic decoder being a key addition. This component mirrors the structure of the density/color decoder and is designed to generate the semantic field in 3D space. For memory efficient training, we incorporate LoRA [16, 41] structures into all linear layers inside both self-attention and cross-attention blocks.
Multiple semantic level supervision. Current LRMs typically rely solely on 2D supervision, which limits their ability to generate information about objects’ internal structures under occlusion; 3D supervision would be effective but often too resource-intensive. To address this, we propose an effective supervision that jointly learns semantics and colors, enabling the acquisition of a 3D semantic field and internal character information using only 2D supervision.
Stage 1: Training on NeRF with Single-layer Semantics. In this initial stage, we train on the triplane NeRF representation. We initialize the model with the pre-trained InstantNeRF, training the newly added LoRA in all attention blocks’ linear layers and the newly introduced semantic decoder. We train it under the image, mask and semantic loss:
@@
Multi-view Comparisons in 2D SyncDreamer [30] 0.870 0.183 0.223 0.864 0.845 0.217 0.328 0.839
@@
Character Comparisons in 3D Magic123 [42] 0.886 0.142 0.192 0.887 0.849 0.197 0.256 0.862
@@
Table 1: Quantitative comparison of A-pose and arbitrary pose inputs for 2D multi-view generation and 3D character generation.
where $\mbox{where }\ T_{P,i}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}\sum_{s\in P}p_{s,j}),$ (8)
The loss function is defined as:
Stage 3: Training on Mesh with Multi-layer Semantics. We switch to mesh representation [49] for efficient high-resolution training. We then extract the equivalent SDF via:
$f_{i,P}=\max(f_{i},(\max_{s\notin P}p_{i,s}-\max_{s\in P}p_{i,s})),$ (10)
$\mathcal{L}_{3}$ $=\mathcal{L}_{2}+\lambda_{\text{normal}}\sum_{k}M^{gt}_{P}\otimes\left(1-\hat{N}_{P,k}\cdot N_{P,k}^{gt}\right)$
$+\lambda_{\text{depth}}\sum_{k}M^{gt}_{P}\otimes\left\|\hat{D}_{P,k}-D_{P,k}^{gt}\right\|_{1}+\lambda_{\text{dev}}\mathcal{L}_{\text{dev}}$ (11)
where $\hat{D}_{P,k}$, $\hat{N}_{P,k}$, denotes the rendered depth and normal; $D_{P,k}^{gt}$, $N_{P,k}^{gt}$ and $M_{P}^{gt}$ denote the ground truth depth, normal, and mask of the $k$-th view under semantic set $P$, respectively; $\mathcal{L}_{\text{dev}}$ denotes the deviation loss of FlexiCubes.
3.3 Multi-layer Refinement
Given the limitations in geometric and texture detail achievable by large reconstruction models, further optimization of the mesh post-reconstruction is necessary. Recent methods [62, 27] utilizing high-resolution normal maps for mesh optimization have shown promising results, albeit primarily for holistic meshes. We propose an iterative optimization mechanism for multi-layer mesh refinement.
To ensure thorough optimization at each level, we employ a staged approach following our S-LRM: Initially, we extract different parts of the mesh by specifying various semantics and optimize only the base minimal-clothed human model; upon completion, we overlay the clothing mesh, fixing the base and optimizing solely the clothing component; finally, we add the hair mesh, fixing the previous two layers and optimizing only the hair. The optimization process is guided by the multi-view normal maps generated earlier through diffusion models, each optimization step involves a differentiable rendering of the mesh to compute losses and gradients, followed by vertex adjustments based on these gradients, and re-mesh operations including edge collapse, split and flips. The loss function is defined as follows:
4 Experiments
@@ 4.1 Anime3D++ Dataset
We develop the Anime3D++ Dataset to meet the quality and decomposability requirements based on the original Anime3D [39] Dataset. We initially gathered around 14,000 3D anime character models from VRoid-Hub, then filtered these down to 10,811 high-quality models, standardized in A-pose with arms angled 45° downward. Renderings include RGB images, depth, normal, and semantic maps for layered supervision. Each character is rendered in full, only a base minimal-clothed human model and a base human model plus clothing. Please refer to the supplementary materials (Sec. E) for more details.
@@ 4.2 Results and Comparisons
User Study. To comprehensively evaluate our method’s performance, we conducted a user study involving 28 volunteers. The study utilized 16 randomly selected 3D meshes and their corresponding reference images, drawn from both web-collected images and the Anime3D++ test split. Participants were asked to compare the randomly shuffled results of four approaches based on overall quality, fidelity, geometric quality, and texture quality. For each comparison group, participants identified the best result in each dimension. Tab. 2 shows that our method achieved superior performance across all dimensions, demonstrating its effectiveness in generating high-quality 3D characters.
@@
Ablation Study. We further conduct ablation studies on character decomposition and multi-layer refinement. Please see supplementary material (Sec. A) for details.
4.3 Applications
Figure 7: Our pipeline enables diverse 3D editing using only text prompts, masks, and in-painting diffusion in the 2D domain.
Our method’s capability to generate decomposed characters in A-pose not only facilitates simpler rigging but also enables 3D editing through user-specified prompts and in-painting masks in the 2D domain. As illustrated in Fig. 7, our framework allows for effortless character customization, including outfit changes and hairstyle modifications. The process requires minimal user input: the original reference image or a front-view capture, a roughly sketched in-painting mask, and a text prompt. Utilizing existing in-painting models (HD-Painter [34] in this case), we achieve the desired 2D edits. Subsequently, these 2D modifications are translated into the 3D domain through our reconstruction and optimization pipeline. During the multi-layer refinement stage, the relevant mesh components are selectively replaced with the edited versions while preserving the unmodified parts. This bridges the gap between 2D image editing and 3D character customization and streamlines the 3D character modification process. We further provide animation comparisons in supplementary material (Sec. B).
5 Conclusion
Acknowledgement. This work was partially supported by Beijing Science and Technology plan project (Z231100005923029), Beijing Natural Science Foundation (L247007), and the Natural Science Foundation of China (62461160309, 62332019).
Supplementary Material
Appendix A Ablation Study
@@
Multi-layer Refinement. We further illustrate the distinction between the direct output of our S-LRM and the results after multi-layer refinement in Fig. 9. The pre-optimization results demonstrate that our S-LRM successfully decomposes various mesh components with correct geometry and shape, validating the capabilities of our S-LRM. However, the precision is limited due to the inherent characteristics of FlexiCubes [49] and memory constraints. Post-refinement, we observe a substantial enhancement in precision while maintaining the overall structure. This improvement underscores the effectiveness of our multi-layer refinement process in preserving the structure of the decomposed components while significantly elevating the geometric accuracy and overall quality of the reconstructed character.
Appendix B More Applications